OpenStudy (anonymous):

For which of the following functions is defined for all real numbers? (choices listed down below)

OpenStudy (anonymous):

A) \[f(x) = \sqrt{3x-2}+ 3\] B) \[f(x) =x^2-16\div x^2-25 \] C) f(x) = \[1\div \sqrt[3]{x}\] D) f(x) = x^1/2 E) f(x) = \[1\div x^2 +4\]

OpenStudy (anonymous):

I wanna say e let me check

OpenStudy (anonymous):

ok, and i thought so too. i just want to be sure. the problems above all look too complicated >.<

OpenStudy (anonymous):

you'll get use to them :D

OpenStudy (anonymous):

you would think by now i should be XD especially after all the math i have done lately XD

OpenStudy (anonymous):

something new is always thrown at you :) dont worry

OpenStudy (anonymous):

to be honest, i have no idea how to go about this problem... cant all of these be real numbers...? @_@

OpenStudy (anonymous):

it means a domain of infinity

OpenStudy (anonymous):

for all rreall numbers

OpenStudy (anonymous):

the correct answer is d.

OpenStudy (anonymous):

oh really? D: would you care to explain? i would love to know

OpenStudy (anonymous):

it's not d at all x^1/2 is sqrt(x) -1 or any negative is not a real number

OpenStudy (anonymous):

it is not D so don't fret about it. you were correct with E

OpenStudy (anonymous):

ooh unles she meant x^1/2 not x^(1/2) :P

OpenStudy (anonymous):

oh thank goodness!! but i was still am confused as to why. i just guessed

OpenStudy (anonymous):

for \(\frac{1}{x^2+4}\) you have to make sure the denominator is not zero but \(x^2+4\geq 4\) for any \(x\) so it is never zero and therefore the domain would be all real numbers

OpenStudy (anonymous):

1/2 is supposed to be a fraction as an exponent btw

OpenStudy (anonymous):

oh wow, that makes much more sense. thanks a bunch satellite73!

OpenStudy (anonymous):

btw when in exponent form put the fraction in parenthesis we humans know what you mean but not a computer :P

OpenStudy (anonymous):

ah, alright. i need to make sure those technicalities are correct XD thank you too

OpenStudy (anonymous):